# Ocean-E2E: Hybrid Physics-Based and Data-Driven Global Forecasting of Extreme Marine Heatwaves with End-to-End Neural Assimilation

### v2

@@ Abstract
This work focuses on the end-to-end forecast of global extreme marine heatwaves (MHWs), which are unusually warm sea surface temperature events with profound impacts on marine ecosystems. Accurate prediction of extreme MHWs has significant scientific and financial worth. However, existing methods still have certain limitations, especially in the most extreme MHWs. In this study, to address these issues, based on the physical nature of MHWs, we created a novel hybrid data-driven and numerical MHWs forecast framework Ocean-E2E, which is capable of 40-day accurate MHW forecasting with end-to-end data assimilation. Our framework significantly improves the forecast ability of extreme MHWs by explicitly modeling the effect of oceanic mesoscale advection and air-sea interaction based on a differentiable dynamic kernel. Furthermore, Ocean-E2E is capable of end-to-end MHWs forecast and regional high-resolution prediction using neural data assimilation approaches, allowing our framework to operate completely independently of numerical models while demonstrating high assimilation stability and accuracy, outperforming the current state-of-the-art ocean numerical forecasting-assimilation models. Experimental results show that the proposed framework performs excellently on global-to-regional scales and short-to-long-term forecasts, especially in those most extreme MHWs. Overall, our model provides a framework for forecasting and understanding MHWs and other climate extremes. Our codes are available at https://github.com/ChiyodaMomo01/Ocean-E2E.
1 Introduction
@@
Marine heatwaves (MHWs) are abnormally warm seawater events that significantly damage marine ecosystems oliver2021marine ; pearce2011marine . In other words, MHWs are extreme Sea Surface Temperature Anomaly (SSTA) events. MHWs, particularly extreme ones, can cause coral bleaching hughes2017global ; hughes2018spatial and widespread mortality of marine organisms garrabou2009mass ; thomson2015extreme . As a result, precise forecasting of extreme MHWs has significant scientific and economic implications. For example, synoptic scale MHWs forecasting can help sea food production and management planning, such as feed cycles, at 1-7 day timescales, while subseasonal to seasonal forecasting can further support proactive decision-making for the blue economy hobday2016seasonal ; malick2020environmentally ; mills2017forecasting ; payne2022skilful . In this study, we will mainly focus on the subseasonal-to-seasonal forecast (i.e., 1-40 day).
Traditional MHW forecasting predominantly relies on numerical models solving oceanic primitive equations, which generally bifurcate into two categories: seasonal forecasts characterizing large-scale patterns of MHW onset and intensification jacox2022global ; brodie2023ecological , and sub-seasonal predictions resolving finer-scale variability with reduced lead times benthuysen2021subseasonal ; yu2024assessing . While these physics-based approaches achieve operational utility, they demand substantial computational resources xiong2023ai ; wang2024xihe and exhibit limited skill in predicting extreme MHW events jacox2022global . Recent advances in deep learning have demonstrated potential for efficient global ocean forecasting bi2022pangu ; chen2023fengwu ; kurth2023fourcastnet , with emerging applications to MHW prediction lin2023rossby ; sun2023artificial . However, existing implementations face dual limitations: 1) insufficient encoding of MHW-specific physics—particularly air-sea coupling mechanisms driving extreme events and mesoscale eddy dynamics governing local intensification—resulting in non-physical smoothing of SST gradients during autoregressive prediction giamalaki2022assessing ; 2) dependence on numerical model initialization that prevent operational end-to-end forecasting from raw observational inputs xiong2023ai . These architectural constraints—shared across mainstream deep learning paradigms—fundamentally limit their capacity to preserve extreme SST anomalies and long-term physical consistency, as error accumulation from neural networks and initialization biases during long-term rollouts wang2024xihe .
@@
By addressing the two shortcomings of the data-driven methods mentioned above, in this study, we designed Ocean-E2E, a global data-driven framework that significantly improves the forecasts of extreme MHWs inspired by numerical models. The contribution of this paper can be summarized as follows: (1) Global MHWs Forecasting Framework. We propose a hybrid physics-based data-driven method that supports both global scale and regional high-resolution forecasting, achieving high-accuracy results for extreme MHWs forecasts. (2) Global MHWs Neural Assmilation Framework. We designed an end-to-end MHWs neural system that, when combined with the forecasting model, can directly obtain the complete initial field and forecasting results of MHWs from observation data. (3) Long-Term Forecast. By combining phyiscal laws into our framework, our model has achieved state-of-the-art results in long-term (subseasonal-to-seasonal) forecasting with higher physical consistency.
The method most similar to us is ClimODE verma2024climode . Similarly, it also uses advection partial differential equations to design the forecasting framework. However, our work has many differences and improvements compared to ClimODE: (1) Differences in theoretical framework. From the perspective of the theoretical framework, in their work, the ‘speed’ of the flow field is essentially an abstract mathematical definition, whose core function is to maintain the conservation of passive tracers. In our work, the flow field speed strictly corresponds to the realistic ocean flow field and models the additional transport effect of the flow field on passive tracers at sub-grid scales through mesoscale ‘bolus’ velocity. (2) Differences in model design. ClimODE uses an emission model to simulate the unclosed and uncertain portions of the real system, while our model further relates this part to the ocean-atmosphere interaction and is designed based on physical theory to form the structure of the source term. (3) Differences in application. ClimODE requires solving the initial velocity field as the initial condition for the model, which involves solving a physics-informed neural network (PINNs) for each training sample. This makes the ClimODE framework difficult to scale up. What’s more, our experiments (Section 4.3) show that our method achieves greater stability and accuracy in MHWs forecasting compared to ClimODE.
@@
2 Problem Setup
@@
In this study, the problem of the end-to-end MHWs forecast is essentially forecasting SSTA, which can be separated into two parts. First, estimating the current state of SSTA $T_{0}\in\mathbb{R}^{H\times W}$ (a global two-dimensional field, where $H$ and $W$ correspond to latitude and longitude) based on the observation $T_{0}^{obs}$ and the background field $T_{0}^{bg}$. Second, predicting the future state of mixed layer SSTA $T_{t}$ based on the current state $T_{0}$ under atmospheric forcing $A_{0:t}\in\mathbb{R}^{T\times H\times W}$. Here the underscore $0:t$ denotes the atmosphere state from time 00 to $t$. Formally speaking, we are going to model the conditional distribution $p(T_{t}|T_{0}^{obs},T_{0}^{bg},A_{0:t},)$ in two stages:
$p(T_{t}|T_{0}^{obs},T_{0}^{bg},A_{0:t})=\underbrace{p^{f}(T_{t}|T_{0},A_{0:t})}_{\text{Forecast model}}\cdot\underbrace{p^{a}(T_{0}|T_{0}^{obs},T_{0}^{bg})}_{\text{Assimilation model}}$ (1)
@@
where $p^{f}$ and $p^{a}$ denote forecast and assimilation process, respectively. In our end-to-end framework, we model these two processes through two specifically designed neural network $\phi^{f}_{\theta}$ and $\phi^{a}_{\theta}$:
$T_{t}=\phi^{f}_{\theta}(T_{0},A_{0:t})$ (2)
$T_{0}=\phi^{a}_{\theta}(T_{0}^{obs},T_{0}^{bg})$ (3)
In summary, our end-to-end framework aims to forecast the furture state of MHWs $T_{t}$ based on the sparse observation $T_{0}^{obs}$ and the initial background field $T_{0}^{bg}$ under the forcing of atmosphere boundary conditions $A_{0:t}$. The parameters of these two neural networks $\theta_{f},\theta_{a}$ are optimized separately by minizing the mean square error (MSE) between the model’s output and the groundtruth:
$\theta_{f}={\rm argmin}\{\frac{1}{N}\sum_{i=0}^{N}(T_{t}^{(i)}-\phi^{f}_{\theta_{f}}(T_{0}^{(i)},A_{0:t}^{(i)}))^{2}\}$ (4)
$\theta_{a}={\rm argmin}\{\frac{1}{N}\sum_{i=0}^{N}(T_{t}^{(i)}-\phi^{a}_{\theta_{a}}(T_{0}^{(i)},A_{0:t}^{(i)}))^{2}\}$ (5)
3 Methods
3.1 Motivation and Framework Overview
@@
Our framework is summarized in Figure 1. As established in the problem formulation, forecasting MHWs fundamentally reduces to predicting extreme sea surface temperature anomalies (SSTA). To effectively embed physical principles into our modeling framework, we must rigorously account for the governing laws of SSTA evolution. We approach this challenge through a generalized geophysical fluid dynamics perspective, wherein seawater temperature constitutes a passive tracer $C$ advected by oceanic flow fields. As shown in Figure 2, the spatiotemporal evolution of $C$ is governed by two primary mechanisms: 1) advective transport by background currents that redistribute thermal energy, 2) mixing and external sink/source encompassing convective mixing and external forcing through surface boundary interactions. Formally speaking, the governing equation of $C$ can be written as:
$\frac{\partial C}{\partial t}+{\bf u}\nabla C+w\frac{\partial C}{\partial z}=\dot{\rm S}+\nabla\cdot(\kappa\nabla C)$ (6)
Where $t$ is time, $({\bf u},w)$ are horizontal and vertical flow velocity, $\nabla$ is the horizontal derivation, $\dot{\rm S}$ is the sink/source term and $\kappa$ is a $2\times 2$ matrix corresponding to horizontal anistropic diffusion coeffiecient smith2004anisotropic . By integrating the above equation, the state of $C$ at time $t$ can be expressed in the form of the initial condition $C_{0}$:
$C_{t}=\int_{0}^{t}-{\bf u}\nabla C+(-w\frac{\partial C}{\partial z}+\dot{\rm S}+\nabla\cdot(\kappa\nabla C))dt+C_{0}$ (7)
@@
Where $-{\bf u}\nabla C$ represents the advective transport while $-w\frac{\partial C}{\partial z}+\dot{\rm S}+\nabla\cdot(\kappa\nabla C)$ corresponds to mixing and external sink/source. Next, we will combine physical theory and deep learning methods to simulate these two mechanisms respectively.
Advective transport. The key factor of advective transport is the ocean surface current ${\bf u}$, where the flow field at the smaller scale (mesoscale) plays an important role in the distribution of MHWs and other oceanic passive tracers. However, in most models and data, due to limited resolution, the model cannot distinguish these small-scale processes, so the advection processes caused by the flow field below the model’s finest grid (i.e., subgrid processes):
$\overline{{\bf u}\nabla C}=\overline{{\bf u}}\overline{\nabla C}+\overline{{\bf u}^{\prime}{\nabla C}^{\prime}}$ (8)
are not fully simulated. Here the overline $\overline{\cdot}$ represents the downsampling operator. In other words, the variables with overline denote the low-resolution data. In the traditional numerical model, Gent and McWilliams gent1990isopycnal found that this subgrid process can be fitted by ‘bolus velocity’ (also called GM90 parameterization):
$\overline{{\bf u}^{\prime}{\nabla C}^{\prime}}\sim{\bf u}_{gm}\overline{\nabla C}$ (9)
According to the GM90 parameterization, these bolus velocity can be expressed asgent1990isopycnal :
@@
where $\rho=\rho(T)$ is the density. It should be noticed that this expression requires subsurface ocean data, which is often extremely difficult to obtain. What’s more, the choice of coefficient $\kappa$ requires empirical formula which is always inaccurate. Alternatively, we seek for a neural network ${\bf u}_{\theta}({\bf u},C)={\bf u}_{gm}$ to approximate the bolus velocity and the subgrid advection effect. In summary, the advection term can be expressed as:
${\bf u}\nabla C=({\bf u}_{\theta}({\bf u},C)+{\bf u})\nabla C$ (11)
Mixing and external sink/source. This part consists of three components: verical mixing, external source, and dissipation. Vertical mixing and dissipation depend more on the instrinic variations of the ocean while the external source $\dot{\rm S}$ is strongly related to the air-sea interaction along the ocean’s surface. To be specific, for SSTA, the source term consists of four parts:
@@
where $Q_{{\rm SW}},Q_{{\rm LW}},Q_{{\rm sens}},Q_{{\rm lat}}$ is the shortwave radiation, longwave thermal radiation, sensible heat flux, and latent heat flux respectively. Of the four air-sea heat fluxes, the sensible and latent flux $Q_{{\rm sens}}^{\prime}+Q_{{\rm lat}}^{\prime}$ determines the majority of the SSTA’s variations holbrook2019global . According to the bulk formula, these two terms are significantly related to surface wind speed $U_{10},V_{10}$, near-surface temperature of the atmosphere $T_{2m}$, and surface specific humidity $q_{a}$. Based on the observation above, we assume that mixing and external sink/source can be well approximated by the surface variables of the atmosphere ${\bf A}=(U_{10},V_{10},T_{2m},q_{a})$ and $C$ itself. Specifically,
$\dot{\rm S}_{\theta}({\bf A},C)=-w\frac{\partial C}{\partial z}+\dot{\rm S}+\nabla\cdot(\kappa\nabla C)$ (13)
In summary, the disscusions above illustrate that the governing processes of SSTA can be efficiently modelled by neural networks. In the following part, we will demonstrate how to incorporate this idea into a hybrid physics-AI model. Specifically, for the bolus velocity approximation ${\bf u}_{\theta}$, we employ a U-Net like architecture. The core of this network, particularly in its deeper layers or bottleneck, utilizes GABlocks (composed of GASubBlock units which integrate SpatialAttention and Mlp modules) to capture multi-scale spatial dependencies from the input ocean current ${\bf u}$ and SSTA $C$. This structure allows the model to learn complex, non-local relationships crucial for representing subgrid advection effects. For the mixing and external sink/source term $\dot{\rm S}_{\theta}$, we utilize a network primarily built upon a stack of GABlocks. This network takes atmospheric variables ${\bf A}$ and the current SSTA $C$ as input, enabling it to effectively model the intricate air-sea interactions and other relevant source/sink processes. The GABlock, with its SpatialAttention mechanism (itself using depth-wise and dilated convolutions as seen in AttentionModule) and Mlp (Multilayer Perceptron with depth-wise convolution, DWConv), is designed to efficiently process and integrate these diverse spatio-temporal inputs for both ${\bf u}_{\theta}$ and $\dot{\rm S}_{\theta}$. The details structure of ${\bf u}_{\theta}$ and $\dot{\rm S}_{\theta}$ can be found in Appendix D.
3.2 Hybrid physics-AI forecast model
As shown in the section above, the furture state of $C$ at time $t$ can be modelled as:
$C_{t}=\int_{0}^{t}-({\bf u}(t)+{\bf u}_{\theta}({\bf u}(t),C)\nabla C+\dot{\rm S}_{\theta}({\bf A}(t),C)dt+C_{0}$ (14)
This indicates that if we are given the values of the ocean currents and atmospheric boundary conditions between the time interval $[0,t]$, we can predict $C_{t}$ in a hybrid physics-AI way.
In our study, the ocean current is obtained from surface geostrophic current inferred from satellite observations. The temporal evolution of this geostrophic current is mainly driven by the ocean instrinic mesoscale dynamics, which can be rigorously described by the 1.5-layer nonlinear RGO model capotondi2001rossby . In other words, the evolution of ${\bf u}$ can be modelled based on an auto-regressive manner:
$\frac{\partial{\bf u}}{\partial t}=\mathcal{M}_{\theta}({\bf u})$ (15)
@@ 3.2 Hybrid physics-AI forecast model
What’s more, the atmosphere variables between time interval $[0,t]$ is also acquired through a large AI-driven weather forecast model gao2025oneforecast :
@@
Here $\mathcal{M}_{\theta},\mathcal{N}_{\theta}$ are two pretrained neural networks. Details of these neural networks can be found in Appendix D. In all, the forecast model can be expressed as:
@@
Where $\hat{C_{t}}$ is the groundtruth data of $C_{t}$. In practice, by separating the time interval $[0,t]$ into $N$ subintervals $[i\Delta t,(i+1)\Delta t],1\leq i\leq N-1$, equation (18) is approximated via the second-order Runge-Kutta method:
$\tiny C_{(i+1)\Delta t}^{\star}$ $=C_{i\Delta t}+\left[-\left({\bf u}_{i\Delta t}+{\bf u}_{\theta}({\bf u}_{i\Delta t},C_{i\Delta t})\right)\nabla C_{i\Delta t}+\dot{\rm S}_{\theta}(A_{i\Delta t:(i+1)\Delta t},C_{i\Delta t})\right]\Delta t$ (19)
$C_{(i+1)\Delta t}$ $=C_{i\Delta t}+\frac{1}{2}\Bigg{\{}\left[-\left({\bf u}_{i\Delta t}+{\bf u}_{\theta}({\bf u}_{i\Delta t},C_{i\Delta t})\right)\nabla C_{i\Delta t}+\dot{\rm S}_{\theta}(A_{i\Delta t:(i+1)\Delta t},C_{i\Delta t})\right]\Delta t$
$\quad+\left[-\left({\bf u}_{(i+1)\Delta t}+{\bf u}_{\theta}({\bf u}_{(i+1)\Delta t},C_{(i+1)\Delta t}^{\star})\right)\nabla C_{(i+1)\Delta t}^{\star}+\dot{\rm S}_{\theta}(A_{(i+1)\Delta t:(i+2)\Delta t},C_{(i+1)\Delta t}^{\star})\right]\Bigg{\}}$ (20)
3.3 End-to-end Neural Data Assimilation

### v3

# Ocean-E2E: Hybrid Physics-Based and Data-Driven Global Forecasting of Marine Heatwaves with End-to-End Neural Assimilation

@@ Abstract
This work focuses on the end-to-end forecast of global extreme marine heatwaves (MHWs), which are unusually warm sea surface temperature events with profound impacts on marine ecosystems. Accurate prediction of extreme MHWs has significant scientific and financial worth. However, existing methods still have certain limitations in forecasting general patterns and extreme events. In this study, to address these issues, based on the physical nature of MHWs, we created a novel hybrid data-driven and numerical MHWs forecast framework Ocean-E2E, which is capable of 40-day accurate MHW forecasting with end-to-end data assimilation. Our framework significantly improves the forecast ability of MHWs by explicitly modeling the effect of oceanic mesoscale advection and air-sea interaction based on a dynamic kernel. Furthermore, Ocean-E2E is capable of end-to-end MHWs forecast and regional high-resolution prediction, allowing our framework to operate completely independently of numerical models while outperforming the current state-of-the-art ocean numerical/AI forecasting-assimilation models. Experimental results show that the proposed framework performs excellently on global-to-regional scales and short-to-long-term forecasts, especially in those most extreme MHWs. Overall, our model provides a framework for forecasting and understanding MHWs and other climate extremes. Our codes are available at https://github.com/ChiyodaMomo01/Ocean-E2E.
Introduction
@@
Marine heatwaves (MHWs) are abnormally warm seawater events that significantly damage marine ecosystems (Oliver et al. 2021; Pearce et al. 2011). In other words, MHWs are extreme Sea Surface Temperature Anomaly (SSTA) events. MHWs, particularly extreme ones, can cause coral bleaching (Hughes et al. 2017, 2018) and widespread mortality of marine organisms (Garrabou et al. 2009; Thomson et al. 2015). As a result, precise forecasting of extreme MHWs has significant scientific and economic implications. For example, synoptic scale MHWs forecasting can help seafood production and management planning, such as feed cycles, at 1-7 day timescales, while subseasonal to seasonal forecasting can further support proactive decision-making for the blue economy (Hobday et al. 2016; Malick et al. 2020; Mills, Pershing, and Hernández 2017; Payne et al. 2022). In this study, we will mainly focus on the subseasonal-to-seasonal forecast (i.e., 1-40 days).
Traditional MHW forecasting has primarily relied on two paradigms: physics-based numerical models and emerging data-driven approaches. Numerical models solve oceanic primitive equations to produce seasonal forecasts that capture large-scale patterns of MHW onset and intensification (Jacox et al. 2022; Brodie et al. 2023), as well as sub-seasonal predictions that resolve finer-scale variability over shorter lead times (Benthuysen et al. 2021; Yu et al. 2024). In parallel, recent advances in deep learning have shown promise for efficient global ocean forecasting (Bi et al. 2022; Chen et al. 2023b; Kurth et al. 2023), with extensions to MHW prediction (Lin et al. 2023; Sun et al. 2023). Despite their operational value, both paradigms suffer from inherent limitations:
1) Numerical models: These approaches demand substantial computational resources (Xiong et al. 2023; Wang et al. 2024). Moreover, they often parameterize key physical processes—such as air-sea coupling and mesoscale eddy dynamics—through empirical formulations, leading to systematic biases and constrained skill in predicting the general patterns of MHWs (Jacox et al. 2022; Giamalaki, Beaulieu, and Prochaska 2022).
2) Data-driven models: While these methods statistically approximate complex oceanic dynamics, they frequently generate non-physical fields, exhibiting limited accuracy in forecasting extreme events like MHWs due to issues such as error accumulation and smoothing of SST gradients during autoregressive rollouts (Wang et al. 2024). Additionally, they typically depend on initialization from numerical models, precluding end-to-end operation from raw observational inputs (Xiong et al. 2023).
Given these complementary strengths and weaknesses, a natural question arises: why not integrate the advantages of both by combining a relatively simple physical model with AI components, where the physical model handles well-understood dynamics and AI simulates the more complex or uncertain aspects? In this study, we pursue this hybrid strategy to effectively mitigate the aforementioned drawbacks of numerical and data-driven models, enabling our framework to harness the best of both worlds:
1) Enhanced computational efficiency and reduced biases: By leveraging AI to parameterize intricate processes like air-sea interactions and eddy effects, we alleviate the heavy resource demands and empirical inaccuracies of traditional numerical models, improving predictions of MHW patterns.
2) Improved physical realism and end-to-end capability: The incorporation of physical constraints ensures generation of consistent fields, addressing the non-physical artifacts and extreme-event limitations in pure data-driven methods, while enabling direct assimilation from observations without reliance on external initializations.
@@
By addressing these integrated shortcomings, we introduce Ocean-E2E, a global hybrid physics-based data-driven framework that significantly enhances forecasts of extreme MHWs, drawing inspiration from numerical models. The contributions of this paper can be summarized as follows: (1) Global/Regional MHWs Forecasting Framework. We propose a hybrid physics-based data-driven method that supports both global scale and regional high-resolution forecasting, achieving high-accuracy results for extreme MHWs forecasts. (2) Global MHWs Neural Assimilation Framework. We designed an end-to-end MHWs neural system that, when combined with the forecasting model, can directly obtain the complete initial field and forecasting results of MHWs from observation data. (3) Physics-consistent Forecast. By combining physical laws into our framework, our model has achieved state-of-the-art results in long-term and regional high-resolution forecasting with higher physical consistency.
@@
Related Work
Numerical Model based MHWs Forecasting
Traditional MHWs forecasting predominantly relies on numerical models solving oceanic primitive equations (Jacox et al. 2022; Brodie et al. 2023; Benthuysen et al. 2021; Yu et al. 2024). While physics-based models demonstrate reasonable forecast skill, they face two principal limitations: prohibitive computational costs that scale with model resolution (Xiong et al. 2023; Wang et al. 2024), and systematic underestimation of extreme MHW magnitudes (Jacox et al. 2022).
Deep Learning based MHWs Forecasting
Machine learning (ML) techniques have emerged as promising alternatives for ocean MHWs modeling, offering orders-of-magnitude speedup over conventional methods (Liu and Ma 2024; Hao et al. 2025). For example, building on advances in atmospheric foundation models (Pathak et al. 2022; Bi et al. 2023), the community has recently proposed global ocean foundation models such as AI-GOMS (Xiong et al. 2023), Xihe (Wang et al. 2024) and WenHai (Cui et al. 2025). However, to our knowledge, there is no AI-based end-to-end MHWs global MHWs forecast model.
Methods
Framework Overview
@@ Methods
In this study, the problem of the end-to-end MHWs forecast is essentially forecasting SSTA, which can be separated into two parts (shown in Figure 1). First, estimating the current state of SSTA $C_{0}\in\mathbb{R}^{H\times W}$ (a global two-dimensional field, where $H$ and $W$ correspond to latitude and longitude) based on the observation $C_{0}^{obs}$ and the background field $C_{0}^{bg}$. Second, predicting the future state of mixed layer SSTA $C_{t}$ based on the current state $C_{0}$ under atmospheric forcing $A_{0:t}\in\mathbb{R}^{C\times H\times W}$. Here $A_{0:t}$ denotes the atmosphere state from time 0 to $t$. Formally speaking, we are going to model the conditional distribution $p(C_{t}|C_{0}^{obs},C_{0}^{bg},A_{0:t},)$ in two stages:
$p(C_{t}|C_{0}^{obs},C_{0}^{bg},A_{0:t})=\underbrace{p^{f}(C_{t}|C_{0},A_{0:t})}_{\text{Forecast model}}\cdot\underbrace{p^{a}(C_{0}|C_{0}^{obs},C_{0}^{bg})}_{\text{Assimilation model}}$ (1)
@@
where $p^{f}$ and $p^{a}$ denote forecast and assimilation process, respectively. In our end-to-end framework, we model these two processes through two specifically designed neural networks $\phi^{f}_{\theta}$ and $\phi^{a}_{\theta}$, which we will discuss in detail in the following sections.
The Design of $\phi^{f}_{\theta}$: Hybrid Physics-AI Model
@@
As established in the problem formulation, forecasting MHWs fundamentally reduces to predicting extreme sea surface temperature anomalies (SSTA). To effectively embed physical principles into our modeling framework, we must rigorously account for the governing laws of SSTA evolution. We approach this challenge through a generalized geophysical fluid dynamics perspective, wherein seawater temperature constitutes a passive tracer advected by oceanic flow fields. Formally speaking, the governing equation of SSTA $C$, together with other surface variables can be written as:
$\frac{\partial C}{\partial t}+({\bf u}_{g}+{\bf u}_{ag})\nabla C+w\frac{\partial C}{\partial z}$ $=\dot{\rm S}$ (2)
$f\mathbf{k}\times\mathbf{u}_{g}$ $=-\nabla(g\eta)$ (3)
[…]
$\frac{\partial\hat{\mathbf{u}}}{\partial t}+\widehat{\mathbf{G}_{u}}+\nabla(g\eta)$ $=0$ (5)
$w|_{z=0}$ $=\eta_{t}$ (6)
Where $t$ is time, $({\bf u},w)=(\mathbf{u}_{g}+\mathbf{u}_{ag},w)$ are horizontal and vertical flow velocity ($\mathbf{u}_{g}$ represents oceanic geostrophic velocity while $\mathbf{u}_{ag}$ is ageostrophic velocity), $\nabla$ is the horizontal derivation, $\dot{\rm S}$ is the sink/source, $\mathbf{k}$ is a unit vector in the vertical direction, $f$ is the Coriolis parameter, $\eta$ is sea surface height, $H$ is ocean depth, $\hat{\cdot}$ is vertical average operator (i.e., $\hat{\mathbf{u}}=\int_{-H}^{0}\mathbf{u}dz$), $g$ is the gravitional accerelation. By integrating equation (1), the state of $C$ at time $t$ can be expressed in the form of the initial condition $C_{0}$:
$\begin{split}C_{t}&=\int_{0}^{t}-{\bf u}_{g}\nabla C+(-{\bf u}_{ag}\nabla C-w\frac{\partial C}{\partial z}+\dot{\rm S})\,dt+C_{0}\\ &\simeq\int_{0}^{t}\underbrace{-{\bf u}_{g}\nabla C}_{\text{advective transport}}+\underbrace{(-{\bf u}_{ag}\nabla C+\dot{\rm S})}_{\text{mixing and external sink/source}}\,dt+C_{0}\end{split}$ (7)
@@
The second equal sign is due to the fact that, according to equation (5), $w|_{z=0}$ is small ($10^{-6}\sim 10^{-5}$ m/s) compared to other terms. Equation (7) implies two key governing mechanisms of MHWs: 1) advective transport by geostrophic currents that redistribute thermal energy, 2) mixing and external sink/source encompassing convective mixing and external ageostrophic forcing through surface boundary interactions. An intuitive description can be found in Figure 2. In the following paragraphs, we will combine physical theory and deep learning methods to simulate these two mechanisms in four steps.
Step 1. Modelling advective transport. The key factor of advective transport is the ocean surface current ${\bf u}_{g}$, where the flow field at the smaller scale (mesoscale) plays an important role in the distribution of MHWs and other oceanic passive tracers. However, in most models and data, due to limited resolution, the model cannot distinguish these small-scale processes, so the advection processes caused by the flow field below the model’s finest grid (i.e., subgrid processes):
$\overline{{\bf u}_{g}\nabla C}=\overline{{\bf u}_{g}}\overline{\nabla C}+\overline{{\bf u}_{g}^{\prime}{\nabla C}^{\prime}}$ (8)
are not fully simulated. Here, the overline $\overline{\cdot}$ represents the downsampling operator. In other words, the variables with an overline denote the low-resolution data. In the traditional numerical model, Gent and McWilliams (Gent and Mcwilliams 1990) found that this subgrid process can be fitted by ‘bolus velocity’ (also called GM90 parameterization):
$\overline{{\bf u}_{g}^{\prime}{\nabla C}^{\prime}}\sim{\bf u}_{gm}\overline{\nabla C}$ (9)
According to the GM90 parameterization, these bolus velocity can be expressed as (Gent and Mcwilliams 1990):
@@
where $\rho=\rho(C)$ is the density. It should be noticed that this expression requires subsurface ocean data, which is often extremely difficult to obtain. What’s more, the choice of coefficient $\kappa$ requires an empirical formula, which is always inaccurate. Alternatively, we seek a neural network ${\bf u}_{\theta}({\bf u}_{g},C)={\bf u}_{gm}$ to approximate the bolus velocity and the subgrid advection effect. In summary, the advection term can be expressed as:
${\bf u}_{g}\nabla C=({\bf u}_{\theta}({\bf u}_{g},C)+{\bf u}_{g})\nabla C$ (11)
Step 2. Modelling mixing and external sink/source. This term is strongly related to the air-sea interaction along the ocean’s surface. To be specific, for SSTA, the source term consists of four parts:
@@
where $Q_{{\rm SW}},Q_{{\rm LW}},Q_{{\rm sens}},Q_{{\rm lat}}$ are the shortwave radiation, longwave thermal radiation, sensible heat flux, and latent heat flux, respectively. Of the four air-sea heat fluxes, the sensible and latent flux $Q_{{\rm sens}}^{\prime}+Q_{{\rm lat}}^{\prime}$ determines the majority of the SSTA’s variations (Holbrook et al. 2019). According to the bulk formula (Cui et al. 2025), these two terms are significantly related to surface wind speed $U_{10},V_{10}$, near-surface temperature of the atmosphere $T_{2m}$, and surface specific humidity $q_{a}$. What’s more, the ageostrophic velocity $\mathbf{u}_{ag}$ is alse driven by surface wind forcing $U_{10}$, $V_{10}$. Based on the observation above, we assume that mixing and external sink/source can be well approximated by the surface variables of the atmosphere ${\bf A}=(U_{10},V_{10},T_{2m},q_{a})$ and $C$ itself. Specifically,
$\dot{\rm S}_{\theta}({\bf A},C)=-{\bf u}_{ag}\nabla C+\dot{\rm S}$ (13)
Since this study primarily focuses on proposing a physics-AI framework, the design and selection of the specific structures for the two neural networks ${\bf u}_{\theta},\dot{\rm S}_{\theta}$ are highly flexible. In this study, we selected a powerful and robust spatiotemporal prediction backbone (Tan et al. 2022) to accomplish this task. We believe that more advanced backbones can further enhance the model’s prediction performance.
Step 3. Treating boundary conditions. Above paragraphs point out that in order to effectively model the key physical mechanism of MHWs, we need the future state of $\mathbf{A}$ and $\mathbf{u}_{g}$ (i.e., boundary conditions). For oceanic geostrophic velocity $\mathbf{u}_{g}$, $\frac{\partial}{\partial t}(3)-\nabla\cdot(H(4))$ implies that
$\frac{\partial^{2}\eta}{\partial t^{2}}+\nabla\cdot(gH\nabla\eta)=\nabla\cdot\widehat{\mathbf{G}_{u}}\simeq\nabla\cdot\widehat{\mathbf{G}_{u_{g}}}\triangleq\mathcal{G}(\eta)$ (14)
which is a self-consistent hyperbolic equation. This indicates that, at least at a relatively short time interval, the evolution of $\eta$, together with geostrophic velocity, can be modelled based on an auto-regressive manner:
$\frac{\partial{\bf u_{g}}}{\partial t}=\mathcal{M}_{\theta}({\bf u_{g}})$ (15)
Here $\mathcal{M}_{\theta}$ is a pretrained neural network. In this study, considering the multi-scale nature of oceanic currents, we design a U-Net-like architecture. The core of this network, particularly in its deeper layers or bottleneck, utilizes GABlocks (composed of GASubBlock units which integrate SpatialAttention and Mlp modules) to capture multi-scale spatial dependencies. This structure allows the model to learn complex, non-local relationships crucial for the oceanic mesoscale motions. More details can be found in Appendix D.
@@
What’s more, the atmosphere variables between time interval $[0,t]$ is also acquired through a large AI-driven weather forecast model (Gao et al. 2025):
@@
Step 4. Combining physics and AI together. As shown in the section above, the furture state of $C$ at time $t$ can be modelled as:
$C_{t}=\int_{0}^{t}-({\bf u}_{g}(t)+{\bf u}_{\theta}({\bf u}_{g}(t),C)\nabla C+\dot{\rm S}_{\theta}({\bf A}(t),C)dt+C_{0}$ (17)
By integrating the boundary conditions mentioned in the above sections, the forecast model can be expressed as:
$\begin{split}C_{t}=&\int_{0}^{t}-\Bigg{[}\int_{0}^{s}\mathcal{M}_{\theta}({\bf u}_{g})ds+{\bf u}_{\theta}(\int_{0}^{s}\mathcal{M}_{\theta}({\bf u}_{g})ds,C)\Bigg{]}\nabla C\\ &+\dot{\rm S}_{\theta}(\int_{0}^{s}\mathcal{N}_{\theta}({\bf A})ds,C)ds\end{split}$ (18)
@@
Where $\hat{C_{t}}$ is the groundtruth data of $C_{t}$. In practice, by separating the time interval $[0,t]$ into $N$ subintervals $i\Delta t,(i+1)\Delta t],1\leq i\leq N-1$, equation (19) is approximated via forward Euler method:
$C_{(i+1)\Delta t}$ $=C_{i\Delta t}+\Big{[}-\big{(}\mathbf{u}_{i\Delta t}+\mathbf{u}_{\theta}(\mathbf{u}_{i\Delta t},C_{i\Delta t})\big{)}\nabla C_{i\Delta t}$
$\quad+\dot{\rm S}_{\theta}(\mathbf{A}_{i\Delta t},\mathbf{A}_{(i+1)\Delta t},C_{i\Delta t})\Big{]}\Delta t,$ (20)
$\mathbf{u}_{(i+1)\Delta t}$ $=\mathcal{M}_{\theta}(\mathbf{u}_{i\Delta t}),$ (21)
$\mathbf{A}_{(i+1)\Delta t}$ $=\mathcal{N}_{\theta}(\mathbf{A}_{i\Delta t}).$ (22)
[…]
The Design of $\phi^{a}_{\theta}$: Neural Data Assimilation